Staff:
Josh, Isabel
Another clear warm day. No interesting weather occurred.
The lenses on the radiometer, METEK lidar, and UVA lidar were cleaned.
The current for both radars was checked around 11:00 AM:
- 449 current - 2.96 A
 - 915 current - 2.45 A
 
Staff:
Josh, Isabel
Another clear warm day. No interesting weather occurred.
The lenses on the radiometer, METEK lidar, and UVA lidar were cleaned.
The current for both radars was checked around 11:00 AM: